# references.bib

### v1

@@ 3.3 Homogenization
Let $A$ be a switch matrix of size $n$ of VC-dimension at most $d$. Then it contains a square $(\alpha,\beta,*)$-matrix with $\alpha\not=\beta$ of size at least $n^{e(4d+1)}$, where $e(\cdot)$ is the function from Theorem 5.
@@
First, we consider the square $(*,0,*)$-submatrix $B$ of $A$ obtained by deleting all even columns of $A$. Let $G$ be the upper graph of $B$. By Observation 2 and Lemma 2, the VC-dimension of $G$ is at most $14d+$ 1. Theorem 5 then implies that $G$ contains a homogeneous subset of vertices $X$ of size at least $n^{e(4d+1)}$. We consider the submatrix $C$ of $A$ with rows $\mathcal{R}=X$, and columns $\mathcal{C}=\{2i-1:i\in X\}$ if $G[X]$ is a clique, resp. $\mathcal{C}=\{2i:i\in X\}$ if $G[X]$ is an independent set. Clearly, $C$ is a square matrix of size at least $n^{e(4d+1)}$.
We claim that the matrix $C$ is an $(\alpha,\beta,*)$-matrix with $\alpha\not=\beta$. If $G[X]$ is a clique, then for each $i,j\in X,i<j$, we have that $A_{i,2j-1}=1$. Moreover, by definition of a switch matrix, we have for all $i\in X$ that $A_{i,2i-1}=0$. Therefore, the matrix $C$ is a $(1,0,*)$-matrix.
If $G[X]$ is an independent set, then for each $i,j\in X,i<j$, we have that $A_{i,2j-1}=0$. Thus, also $A_{i,2j}=0$ by definition of a switch matrix. Moreover, we have for all $i\in X$ that $A_{i,2i}=1$. Therefore, the matrix $C$ is a $(0,1,*)$-matrix. ∎
@@
Let $A$ be a square $(\alpha,\beta,*)$-matrix of size $n$ of VC-dimension at most $d$. Then it contains a square $(\alpha,\beta,\gamma)$-matrix with $\alpha\not=\beta$ of size at least $n^{e(4d+1)}$, where $e(\cdot)$ is the function from Theorem 5.
@@
We proceed similarly as in the proof of Lemma 3, except that we construct the upper graph from the transposition of the matrix $A$, resp. from the the binary complement of the transposition if $A$ is a $(0,1,*)$-matrix (to formally complain with Lemma 2 by having 00 on the diagonal). By applying Theorem 5, we obtain a square $(\alpha,\beta,\gamma)$-submatrix of $A$ with $\alpha\not=\beta$ of size at least $n^{e(4d+1)}$. ∎
@@
By composing Corollary 1 with Lemmas 3 and 4, we obtain an $(\alpha,\beta,\gamma)$-diagonal submatrix with $\alpha\not=\beta$ of polynomial size with the exponent
$\frac{1}{d}\cdot e(4d+1)\cdot e(4d+1),$
@@ 4 Remarks on tame classes
In recent years, much effort has been invested into combinatorial understanding of classical model-theoretic notions such as stability and dependence with the aim to establish dividing lines between tame and wild classes of finite structures [Gajarsky2023][Torunczyk2023][Dreier2024]. These include the model-theoretic notions of (monadic) stability, (monadic) dependence and combinatorial notion of bounded twin-width. For their definitions please see [Tent2012][Simon2015][Dreier2024][Bonnet2021], Figure 1 shows their inclusions. Considered folklore, Theorem 1 plays an important role in this area. For example, it was one of the tools that allowed Dreier, Mählmann, and Toruńczyk to characterize the forbidden subgraphs of monadically dependent graph classes, yielding the first purely combinatorial definition of monadic dependence [Dreier2024].
@@
The coarsest of these notions, dependence, implies bounded VC-dimension. Therefore, our improvement in form of Theorem 2 applies for dependent classes of graphs and for all classes from Figure 1. However, the more restricted classes allow for further improvements of Theorem 2 and better bounds.
In Corollary 1, we used Theorem 4 to bound the growth function of classes of bounded VC-dimension. While Theorem 4 gives a polynomial bound on the growth function of the neighborhood systems, also called neighborhood complexity, it is conjectured that the monadically dependent classes admit an almost linear bound [Dreier2023a][Dreier2024]. That is, that for graphs from the class $\mathcal{C}$, the growth function is bounded by $bn^{1+\varepsilon}$ for every $\varepsilon>0$ so some constant $b=b(\varepsilon,\mathcal{C})$. This is known to be true for monadically stable classes [Dreier2023b]. For classes of bounded twin-width, the situation is even more pleasant as the neighborhood complexity is linear as proved independently by [Bonnet2022] and [Przybyszewski2023], with an essentially tight bound by [Bonnet2024a]. Using these exponents in Lemma 1, we conclude that graphs from classes of bounded twin-width admit a switch submatrix of a linear size, while monadically stable classes admit a polynomial switch submatrix with the exponent $1-\varepsilon$ for every $\varepsilon>0$.
Moreover, if the upper graph of the given matrix belongs to a tame class, we can reach better bounds by substituting Theorem 5 by a more specific statement. It was implicitly proved by Malliaris and Shelah [Malliaris2013, Theorem 3.5] that stable graph classes posses the Erdős-Hajnal property. The bound was further improved in [Malliaris2017, Corollary 4.9] using only the milder assumption that the graph class is edge-stable (that is, only the edge relation is required to be stable; not necessarily all the definable relations). Furthermore, for the more restricted monadically stable graph classes, it was recently proved that they possess the Erdős-Hajnal property with the exponent $1/2-\varepsilon$ for any $\varepsilon>0$ [Braunfeld2025, Corollary 62]. Notably, this says that there is a uniform Erdős-Hajnal exponent for all monadically stable graph classes.
Speaking about stable classes, we should point out that we can further simplify the conclusion of Theorem 2. Indeed, for each stable graph class there is a bound on the size of half-graphs that can appear as subgraphs, which allows us to narrow the conclusion of Theorem 2 that the target subgraph is either a matching or a co-matching.
@@ 5 Acknowledgement
The author would like to thank Patrice Ossona de Mendez for his many suggestions concerning Section 4.
The author is partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC Synergy Grant DYNASNET, grant agreement No 810115), and by Project 24-12591M of the Czech Science Foundation (GAČR).
[…]

### v2

@@ 3.3 Homogenization
Let $A$ be a switch matrix of size $n$ of VC-dimension at most $d$. Then it contains a square $(\alpha,\beta,*)$-matrix with $\alpha\not=\beta$ of size at least $n^{e(4d+3)}$, where $e(\cdot)$ is the function from Theorem 5.
@@
First, we consider the square (,0,)(*,0,*)-submatrix $B$ of $A$ obtained by deleting all even columns of $A$. Let $G$ be the upper graph of $B$. By Observation 2 and Lemma 2, the VC-dimension of $G$ is at most $4d+3$. Theorem 5 then implies that $G$ contains a homogeneous subgraph on a vertex set $X$ of size at least $n^{e(4d+3)}$. We consider the submatrix $C$ of $A$ with rows $\mathcal{R}=X$, and columns $\mathcal{C}=\{2i-1:i\in X\}$ if $G[X]$ is a clique, or $\mathcal{C}=\{2i:i\in X\}$ if $G[X]$ is an independent set. Clearly, $C$ is a square matrix of size at least $n^{e(4d+3)}$.
We claim that the matrix $C$ is an $(\alpha,\beta,*)$-matrix with $\alpha\not=\beta$. If $G[X]$ is a clique, then for each $i,j\in X,i<j$, we have that $A_{i,2j-1}=1$. Moreover, by definition of a switch matrix, we have for all $i\in X$ that $A_{i,2i-1}=0$. Therefore, the matrix $C$ is a (1,0,)(1,0,*)-matrix.
If $G[X]$ is an independent set, then for each $i,j\in X,i<j$, we have that $A_{i,2j-1}=0$. Thus, also $A_{i,2j}=0$ by definition of a switch matrix. Moreover, we have for all $i\in X$ that $A_{i,2i}=1$. Therefore, the matrix $C$ is a (0,1,)(0,1,*)-matrix. ∎
@@
Let $A$ be a square $(\alpha,\beta,*)$-matrix of size $n$ of VC-dimension at most $d$. Then it contains a square $(\alpha,\beta,\gamma)$-matrix of size at least $n^{e(4d+3)}$, where $e(\cdot)$ is the function from Theorem 5.
@@
We proceed similarly as in the proof of Lemma 3, except that we construct the upper graph from the transpose of the matrix $A$, or from the binary complement of the transpose if $A$ is a (0,1,)(0,1,*)-matrix (to formally comply with Lemma 2 by having 0 on the diagonal). By applying Theorem 5, we obtain a square $(\alpha,\beta,\gamma)$-submatrix of $A$ of size at least $n^{e(4d+3)}$. ∎
@@
By composing Corollary 1 with Lemmas 3 and 4, we obtain an inhomogeneous square $(\alpha,\beta,\gamma)$-submatrix with $\alpha\not=\beta$ of polynomial size with the exponent
$\frac{1}{d}\cdot e(4d+3)\cdot e(4d+3),$
@@ 4 Remarks on tame classes
In recent years, much effort has been invested into combinatorial understanding of classical model-theoretic notions such as stability and dependence with the aim to establish dividing lines between tame and wild classes of finite structures [21][35][16]. These include the model-theoretic notions of (monadic and edge-)stability, (monadic) dependence and the combinatorial notion of bounded twin-width. For their definitions please see [34][33][16][8], Figure 1 shows their inclusions. Considered folklore, Theorem 1 plays an important role in this area. For example, it was one of the tools that allowed Dreier, Mählmann, Siebertz, and Toruńczyk to give the first purely combinatorial characterization of monadic stability [15]. Moreover, Dreier, Mählmann, and Toruńczyk were able to describe the forbidden subgraphs of monadically dependent graph classes, yielding the first combinatorial characterization of monadic dependence [16].
@@
All the notions in Figure 1 imply bounded VC-dimension. Therefore, our improvement in form of Theorem 2 applies to all these classes. However, the more restricted notions allow for further improvements of Theorem 2 and better bounds.
In Corollary 1, we used Theorem 4 to bound the growth function of classes of bounded VC-dimension. While Theorem 4 gives a polynomial bound on the growth function of the neighborhood systems, also called neighborhood complexity, it is conjectured that the monadically dependent classes admit an almost linear bound [14][16]. That is, that for graphs from the class $\mathcal{C}$, the growth function is bounded by $cn^{1+\varepsilon}$ for every $\varepsilon>0$ a constant $c=c(\varepsilon,\mathcal{C})$. This is known to be true for monadically stable classes [13]. For classes of bounded twin-width, the situation is even more pleasant as the neighborhood complexity is linear as proved independently by [7] and [29], with an essentially tight bound by [4]. Using these exponents in Lemma 1, we conclude that graphs from classes of bounded twin-width admit a switch submatrix of a linear size, while monadically stable classes admit a polynomial switch submatrix with the exponent $1-\varepsilon$ for every $\varepsilon>0$.
Moreover, if the upper graph of the given matrix belongs to a tame class, we can reach better bounds by substituting Theorem 5 by a more specific statement. This is the case of classes of bounded twin-width, [^2] Similarly as in Lemma 2, large twin-width of $G$ is witnessed by a certain submatrix called mixed-minor [8]. Then $A$ contains a mixed-minor of at least half of the size, and consequently has large twin-width as well. The contrapositive states that the upper graph of a matrix with bounded twin-width has itself bounded twin-width. However, a more direct argument with a better quantitative bound would be desirable. which admit even the strong Erdős-Hajnal property [5, Theorem 20]. It was implicitly proved by Malliaris and Shelah [25, Theorem 3.5] that stable graph classes possess the Erdős-Hajnal property. The bound was further improved in [26, Corollary 4.9] using only the milder assumption of edge-stability (that is, only the edge relation is required to be stable; not necessarily all the definable relations). Furthermore, for the more restricted monadically stable graph classes, it was recently proved that they possess the Erdős-Hajnal property with the exponent $1/2-\varepsilon$ for any $\varepsilon>0$ [9, Corollary 62]. Notably, this says that there is a uniform Erdős-Hajnal exponent for all monadically stable graph classes.
Speaking about (edge-)stable classes, we should point out that we can further simplify the conclusion of Theorem 2. Indeed, for each edge-stable graph class there is (by definition) a bound on the size of half-graphs that can appear as subgraphs, which allows us to narrow the conclusion of Theorem 2 that the target subgraph is either a matching or a co-matching.
@@ 5 Acknowledgement
The author would like to thank Patrice Ossona de Mendez for his many suggestions concerning Section 4. Many thanks also belong to the reviewers whose helpful comments have helped improve the presentation.
The author is partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC Synergy Grant DYNASNET, grant agreement No 810115), by Project 24-12591M of the Czech Science Foundation (GAČR), and by the grant SVV–2025–260822.